OpenStudy (anonymous):

What is the sum of the roots of the polynomialshown below? f(x)=3x^3+12x^2+3x-18

OpenStudy (anonymous):

factor a 3 and u can "see" 1 root

jimthompson5910 (jim_thompson5910):

Hint: The sum of the roots of the polynomial \[a_nx^n+a_{n-1}x^{n-1}+\cdots+a_1x+a_0\] is \[-\frac{a_{n-1}}{a_n}\] See link below for more info http://www.slideshare.net/kliegey524/polynomials-sum-and-product-of-the-roots

OpenStudy (anonymous):

Didn't know that formula, I checked, -6,1,1. Thanks.

jimthompson5910 (jim_thompson5910):

It turns out that the roots are -3,-2, and 1. Add them up to get -4. We can see that \[a_{n-1}=a_{2}=12\] and \[a_n=a_3=3\], so the sum of the roots is \[-\frac{a_{n-1}}{a_n}=-\frac{12}{3}=-4\]

OpenStudy (anonymous):

Have to go to polynomial classes, lol

OpenStudy (anonymous):

lol yea